OpenStudy (trojanpoem):

integrate:

OpenStudy (trojanpoem):

\[\int\limits_{}^{}\frac{ \sin^2x - 4sinxcosx + 3 \cos^2x }{ sinx + cosx } dx\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

maybe try \[\int \frac{1+2\cos^2(x)-4\sin(x)\cos(x)}{\sin(x)+\cos(x)}dx\]?

OpenStudy (trojanpoem):

I tried long division

OpenStudy (misty1212):

no, that doesn't seem to help

Vocaloid (vocaloid):

if i'm not mistaken, can the numerator be factored?

OpenStudy (photon336):

there's got to be a way to simplify the integral

OpenStudy (trojanpoem):

I ended up with sinx - 5 cosx + (8 cos^2x)/(sinx+cosx)

OpenStudy (trojanpoem):

Vocaloid, yeah I fatored it (sinx - 3cosx)(sinx- cosx)

OpenStudy (trojanpoem):

After long division I had to solve this integral cos^2x/(sinx+cosx)

OpenStudy (faiqraees):

That 2cos²x in misty's integral can be converted into cos2x +1

OpenStudy (misty1212):

i tried it with wolfram, solution is pages long and a big fat mess, but sometimes wolfram doesn't know the easy way

OpenStudy (trojanpoem):

Not pages

OpenStudy (trojanpoem):

3cosx - sinx + 2sqrt(2) ln((sqrt(2)-1)cosx +sinx/(sqrt(2)-1)cosx-sinx)) + c

OpenStudy (trojanpoem):

That's the final result.

OpenStudy (faiqraees):

maybe try \[\int \frac{2+cos(2x)-2\sin(2x)}{\sin(x)+\cos(x)}dx\]?

OpenStudy (trojanpoem):

@FaiqRaees , Integrating different angles is a disaster SO HARD.

OpenStudy (bobo-i-bo):

How I went about to solve it, it's a bit convoluted: \[\int\limits \frac{ \sin^2x-4sinxcosx+3\cos^2x }{ sinx + cosx }dx=\] \[\int\limits \frac{ (sinx -cosx)(sinx-3cosx) }{ sinx + cosx }dx\] Using substitution \(u=sinx + cosx \Rightarrow -du =(sinx -cosx) dx \): \[\int\limits \frac{ (sinx -cosx)(sinx-3cosx) }{ sinx + cosx }dx=\] \[\int\limits \frac{ -u+4cosx }{ u } du=-\int\limits 1 du +4 \int\limits \frac{ cosx }{ u } du=\] \[-u +4 \int\limits \frac{cosx}{u} du = -sinx - cosx + 4 \int\limits \frac {cosx(cosx - sinx)}{cosx + sinx}dx\]

OpenStudy (bobo-i-bo):

Now: \[cosx(cosx-sinx)=\cos^2x-cosxsinx=1-sinx^2-cosxsinx=1-sinx(sinx+cosx)\] So: \[\int\limits \frac{cosx(cosx-sinx)}{cosx+sinx}dx= \int\limits \frac{1-sinx(sinx+cosx)}{cosx+sinx}dx=\] \[\int\limits \frac 1 {cosx + sinx} dx - \int\limits sinx dx\] Finally: \[cosx+sinx=\sqrt 2 \sin(x+\frac {\pi} 2)\] So now I should have given you everything necessary to piece it all together and fill in the other bits and pieces. :P

OpenStudy (trojanpoem):

@Bobo-i-bo, Great job mate.